# Task-driven SLAM Benchmarking for Robot Navigation

### v2

@@ Abstract
A critical use case of SLAM for mobile assistive robots is to support localization during a navigation-based task. Current SLAM benchmarks overlook the significance of repeatability (precision), despite its importance in real-world deployments. To address this gap, we propose a task-driven approach to SLAM benchmarking, TaskSLAM-Bench. It employs precision as a key metric, accounts for SLAM’s mapping capabilities, and has easy-to-meet implementation requirements. Simulated and real-world testing scenarios of SLAM methods provide insights into the navigation performance properties of modern visual and LiDAR SLAM solutions. The outcomes show that passive stereo SLAM operates at a level of precision comparable to LiDAR SLAM in typical indoor environments. TaskSLAM-Bench complements existing benchmarks and offers richer assessment of SLAM performance in navigation-focused scenarios. Publicly available code permits in-situ SLAM testing in custom environments with properly equipped robots.
@@ I Introduction
Service robots have increasingly been deployed into home and work environments, as well as other parts of daily life [1, 2, 3, 4, 5, 6, 7]. Their primary goal is to provide repeatable services that enhance productivity and foster more efficient work environments. Central to the operation of these mobile robots is SLAM, whose localization capabilities support navigation. A robust and reliable SLAM system is pivotal for completing tasks that are distributed throughout a given environment. Despite recent advancements in SLAM achieving high accuracy in existing benchmarks and datasets [8, 9, 10, 11], service robots using SLAM for feedback continue to face challenges with robustness such that localization failures require human intervention to restore functionality [12]. These closed-loop failures are not adequately addressed by existing open-loop SLAM benchmarks, which means that performance in the former does not always translate to task-oriented autonomy performance. For service robots, the main concern is reliable navigation to the same location when needed, rather than highly accurate absolute positioning. Additional perception-based modules ensure completion of the service activity near the goal location. Consider that navigation benchmarks focus on robot terminal point performance, where success is defined by the robot reaching the goal region [13, 14]. A robot needs sufficient localization performance to reach the goal region. Of essence is reliable, repeatable performance to ensure successful task execution every time. Repeatability, as measured using precision, is a common metric in robot automation [15, 16, 17, 18]. This paper advocates to augment open-loop SLAM benchmarks with navigation-based SLAM benchmarks to better evaluate operational performance, and describes an easy to implement benchmarking scheme.
Figure 1: Precision vs. Accuracy SLAM performance results without (w/o) and with (w/) a map for EuRoC Machine Hall sequences. Each marker is a single sequence under a specific test mode. The dashed red line marks where precision equals accuracy. All data points being below this line suggest that accuracy under-reports task-repeatability. Metrics and units are defined in §II-B2, with experimental details given in §III-C.
@@
Real-world benchmarking with trajectory estimation error metrics remains mostly open-loop due to the metrics requiring absolute, global ground truth signals. Significant effort is required to obtain these signals. For instance, KITTI [8] uses RTK-GPS for precise vehicle positioning, and HILTI [9] employs a total station (e.g., a highly accurate laser scanner), both of which are costly. Datasets like EuRoC [10] and TUM [19] use motion-capture systems, which are difficult to implement in large-scale indoor, multi-room environments. For larger scale, multi-room environments, methods like PennCOSYVIO [20] use fiducial markers in conjunction with a laborious process involving meticulous placement and measurement of the markers. Sparser ceiling-mounted marker placement with total station surveying and an extra upwards facing camera on the robot can be a more effective approach for indoor settings. Other methods [21, 22] rely on LiDAR-SLAM or multi-sensor fusion results from an offline pose optimization process as ground truth. The proposed method removes the need for costly absolute measurement technology, labor-involved annotation, and offline pose optimization for ground truth.
@@
The Subterranean (SubT) Challenge [23] was a real-world challenge requiring closed-loop implementation (e.g., SLAM and navigation integration) during exploration in underground environments with the objective of localizing targets distributed throughout them. It prioritized target detection and target position in an absolute global frame. The evaluation criteria emphasized giving the true location of the target rather than being able to reliably guide a human or robot back to the target. Resource intensive surveying methods were used to generate the ground truth.
@@
While significant efforts have recently been devoted to rich, multi-sensor fusion capable benchmarking [21, 9], the performance gap between SLAM benchmarking via open-loop, sensor stream replay versus task-oriented performance has motivated other to evaluate SLAM and related autonomy modules through different lenses. Simulated worlds figure prominently due to reproducibility of outcomes, ease of implementation by researchers, and the availability of ground truth information. Embodied AI research, such as the Habitat Navigation Challenge [24], exemplifies the necessity for a task-driven evaluation methodology. The simulation-based challenge emphasizes the need for reliable, long-term navigation capabilities. The potential of localization to support autonomous vehicles has also led to the design of a simulation-based benchmarking scheme for street navigation [25] with success/failure as the main metric. SLAM latency plays a role in its robustness and accuracy, with some relationship found between these two factors for the task of trajectory tracking using SLAM pose estimates as feedback [26, 27]. The simulation-based benchmarking in [26, 27] demonstrate the need to move beyond just accuracy when considering SLAM in the closed-loop. A reasonable next step would be to more explicitly consider navigation.
@@
An additional difference between these benchmarks and the deployment of service robots is that the latter often incorporate an initial mapping stage prior to being tasked. Existing benchmarks overlook this phase by evaluating performance solely through one-time data playback. SLAM maps are pivotal as they mitigate long-term drift and ensure consistency (high precision) across multiple visits to the same locations [28, 29]. Multi-session SLAM evaluation [30] does consider map reuse with accuracy metric for repeated runs to the same locations. However, the study design and variables may consider different performance factors and experimental methodology for LiDAR SLAM [30], and also for visual SLAM multi-map and map reuse open-loop tests [28, 29]. These works point to the need for reproducible and standardized benchmarking schemes for SLAM systems equipped with pre-built maps. Further evidence for this need is given in Fig. 1, where the EuRoC dataset is used in a multi-run manner (see §III-C for more details). The Precision vs Accuracy plots show better precision than accuracy indicating that SLAM evaluation may under-report performance. Including a mapping phase further improves performance, especially position precision. Referencing the performance metrics to robot characteristics (under $x$-axis) permits qualitative assessment of potential goal attainment success by relating precision to what task region the robot may see at its arrival pose.
@@
This paper introduces a SLAM benchmark framework with navigation-in-loop that emphasizes task-relevant metrics such as repeatability (precision) and completion. The benchmarking approach mimics real-world implementations where robots are typically given ample time to create a complete map for improved task execution during deployment. Through experiments on several visual and LiDAR SLAM methods, we identify which methods effectively support navigation tasks. The key contributions are as follows:
@@
  • A low-cost, easy to setup, task-driven SLAM benchmark with precision as the key metric for measuring robot consistency in reaching the same pose over multiple rounds. The method readily scales to large environments.
@@ II-A Preliminaries
  • Robot Pose - $(x,y,\theta)$ Robot position and orientation. $SE(3)$ pose estimates are projected to $SE(2)$.
  • Sequence - $\mathbf{S}$ A set of waypoints the robot is required to sequentially follow, i.e., $\mathbf{S}=\{\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{N}\}$.
  • Waypoint - $\mathbf{w}_{i}=(x_{i},y_{i},\theta_{i})$ indexed by $i$ position and orientation (heading angle) of the robot in the 2D plane.
@@ II-B Task-Driven SLAM Benchmarking Design
Figure 2: Task-driven SLAM benchmarking design. Dashed arrows and cylinders represent optional elements specific to SLAM methods and test configurations. The ||||| | symbol denotes alternative robot sensor configurations.
@@ II-B1 Task Definition
Waypoint Navigation. Waypoints represent task-relevant goal points, as exemplified in the AWS Gazebo hospital model [34] of Fig. 3. They are manually specified based on a floor plan of the environment’s free space. It is assumed that the floor plan reflects the static environment, with no dynamic obstacles or unknown areas. The robot starts from a fixed location and sequentially navigates to each waypoint. Upon reaching a waypoint, the robot’s pose is recorded for evaluation. The procedure is repeated for multiple rounds, terminating if the robot fails to reach any waypoint.
@@ II-B2 Performance Metrics
Accuracy, Completeness, and Precision. Typically, accuracy metrics such as Absolute Trajectory Error (ATE) or Relative Pose Error (RPE) are used to evaluate SLAM performance [19]. Navigation and trajectory tracking evaluation of SLAM employ task-specific metrics [25, 26]. Here, performance evaluation of indoor mobile robot navigation emphasizes waypoint navigation reliability and repeatability, measured by successful goal attainment percentage (completeness) and goal pose precision. Accuracy, in the form of APE, will apply for the simulated scenarios since ground-truth is available.
@@
Task-driven Units - $(D,FoV)$. To emphasize performance from a task-driven perspective, outcomes will also report in unit scales relevant to the robot’s physical characteristics: $D$ the robot’s diameter, and $FoV$ the camera’s horizontal field of view. These units to quantify position and orientation performance are robot-specific.
@@ II-B3 Measuring Completeness and Precision
Waypoint precision measurement in real-world conditions may be obtained in two equivalent manners similar to virtual reality systems: outside-in or inside-out. Outside-in uses a downward facing camera from a suitable height (Fig. 4). Options include overhead/ceiling-mounted, tripod-installed, or otherwise placed to face the ground region the robot is expected to reach. Each camera requires connection to a (low cost) computer or laptop to estimate the robot’s pose upon reaching the waypoint. An AprilTag is affixed to the top of the robot for pose estimation from the camera system (Fig. 4). AprilTag detection validation [35] indicates point error is within 0.5 pixel for the AprilTag C++ library, making it suitable for evaluation purposes. Inside-out operates as in [36] by inverting the setup. An upwards-facing camera is placed on the robot and tags are placed in the environment.
Importantly, waypoint precision does not rely on absolute references. It is assessed independently of other waypoints, and the evaluation is possible in local coordinate systems defined by the cameras observing them. The outside-in approach does not require global calibration of the camera poses (nor relative poses). Each camera/waypoint measurement pair operates independently. The inside-out approach approach does not require a global scan nor globally aligned tags. Lastly, waypoint attainment success is established by the tag being fully visible to the camera, otherwise the robot failed to reach the waypoint’s vicinity.
@@
Figure 4: Real-world experimental setup. Clock-wise from top-left: A Turtlebot2 is equipped with the SLAM sensors and laptop, on top of which is placed an AprilTag. Downward facing cameras (in red box) and computers/laptops are placed in the environment to detect and recover the robot’s pose. The robot waypoints, defined to be under the cameras, should be accurate enough that the robot will be visible when reached. During the programmed tour, the cameras will detect and estimate the robot pose when seen. Precision and Completeness follows from the estimates. The lower figure presents an example of the robot captured at two waypoints (magenta and green) for different SLAM methods. Arrows indicate orientation and black circles indicates position, over multiple runs.
@@ III Experiment Design
The effectiveness of the proposed benchmark framework will be established through a series of controlled experiments. Simulation experiments assess SLAM performance using all metrics: Accuracy, Precision, and Completeness. Plots will include dual axes with standard units and robot-relative units to link performance directly to the robot’s physical characteristics and task requirements. If a robot is too far from an object, relative to its size, or headed in a bad direction, relative to its visual field of view, then a follow-up visually guided manipulation, inspection/object viewing, or other task would fail.
@@ III-A SLAM Candidates
To demonstrate the flexibility of the benchmark and support comparative analysis, several SLAM methods are tested. They include inertial and robot odometry signals as required or when permitted. SLAM parameters were not tuned for optimal performance across the tests, with default settings used; they are generally effective in most scenarios and offer balanced performance. Our goal is not to identify the best SLAM method, but rather to demonstrate the functionality of the proposed SLAM benchmarking design (TaskSLAM-Bench), and to use the demonstration to assess which SLAM methods are suitable for task-driven purposes. Tested are:
@@ III-A2 Stereo Visual and Visual-Inertial
GF-GG[27] & ORB_SLAM3[28], feature-based stereo visual methods; DSOL[39], a direct sparse odometry system; SVO-Pro[40], a semi-direct method; and MSCKF[41], a filter-based method. ORB_SLAM3 would often fail to initialize when using inertial measurements, even with extra movement prior to testing. It is run as stereo-only.
@@
III-B SLAM-in-the-Loop Navigation Experiments
Testing involves both simulated and real-world environments, whose scenarios are described in Table I. In each scenario, the robot starts from a fixed initial location, typically the origin of the map, and navigates sequentially through the waypoints. The navigation system configuration is fixed across all experiments. Upon reaching each waypoint, the robot pauses for 5 seconds before proceeding to the next one. For an outside-in setup, the computer/laptop at this waypoint, synchronized with NTP (delay $<$ 1 second), saves timestamped image data upon detecting the tag affixed to the robot. The robot’s poses are estimated offline and associated with waypoints by timestamp for performance evaluation.
@@ III-B2 Real World
Real-world testing was performed in an office environment. One test defined a path sequence of $\sim\!45$ meters in length and featuring six waypoints, see Fig. 4. It has short stretches in a corridor, sharp turns, and return loops. Using an outside-in approach, three overhead cameras (Fig. 4) captured the robot for success determination and precision measurements. Another test, with the inside-out approach, defined a travel sequence of about 120 meters in length with twelve waypoints. It has similar properties to the first test but over longer distances.

### v3

@@ Abstract
A critical use case of SLAM for mobile robots is to support localization during task-directed navigation. Current SLAM benchmarks overlook the importance of repeatability (precision) despite its impact on real-world deployments. TaskSLAM-Bench, a task-driven approach to SLAM benchmarking, addresses this gap. It employs precision as a key metric, accounts for SLAM’s mapping capabilities, and has easy-to-meet requirements. Simulated and real-world evaluation of SLAM methods provide insights into the navigation performance of modern visual and LiDAR SLAM solutions. The outcomes show that passive stereo SLAM precision may match that of 2D LiDAR SLAM in indoor environments. TaskSLAM-Bench complements existing benchmarks and offers richer assessment of SLAM performance in navigation-focused scenarios. Publicly available code permits in-situ SLAM testing in custom environments with properly equipped robots.
@@ I Introduction
Service robots have increasingly been deployed into home and work environments, and in public spaces [1, 2, 3, 4, 5, 6, 7]. Their primary goal is to provide repeatable services that enhance productivity and foster more efficient work environments. Central to the operation of these mobile robots is SLAM, whose localization capabilities support navigation. Robust and reliable SLAM is pivotal to completing tasks that are distributed throughout a given environment. Despite recent advancements in SLAM achieving high accuracy in existing benchmarks and datasets [8, 9, 10, 11], service robots using SLAM for feedback continue to face challenges with robustness such that localization failures require human intervention to restore functionality [12]. These closed-loop failures are not adequately addressed by existing open-loop SLAM benchmarks, such that high performance in the former does not always translate to autonomous task-oriented performance. For service robots, the main concern is reliable navigation to the same location when tasked, rather than high accuracy absolute positioning. Additional perception-based modules ensure completion of the service activity near the goal location. Consider that navigation benchmarks focus on robot terminal point performance, where success is defined by the robot reaching the goal region [13, 14]. A robot needs sufficient localization performance to reach the goal region. Of essence is reliable, repeatable performance to ensure successful task execution every time. Repeatability, as measured using precision, is a common metric in robot automation [15, 16, 17, 18]. This paper advocates to complement open-loop SLAM benchmarks with navigation-based SLAM benchmarks to better evaluate operational performance, and describes an easy to implement benchmarking scheme.
[…]
@@
Real-world benchmarking with trajectory estimation error metrics [19] remains mostly open-loop due to metrics requiring absolute, global ground truth signals. Significant effort is required to obtain these signals. For instance, KITTI [8] uses RTK-GPS for precise vehicle positioning, and HILTI [9] uses a total station (e.g., a survey grade laser scanner), both of which are costly. Datasets like EuRoC [10] and TUM [20] use motion-capture systems, which are difficult to implement in large-scale indoor, multi-room environments. In these cases, methods like PennCOSYVIO [21] use fiducial markers in conjunction with a laborious process involving meticulous placement and measurement of the markers. Sparse ceiling-mounted marker placement with total station surveying and an extra upwards facing camera on the robot is effective for indoor settings [22]. Other methods [23, 24] rely on LiDAR-SLAM or multi-sensor fusion results from an offline pose optimization process as ground truth. The proposed method removes the need for absolute measurement technology, labor-involved annotation, and offline pose optimization.
@@
The Subterranean (SubT) Challenge [25] was a real-world challenge requiring closed-loop implementation (e.g., SLAM and navigation integration) during exploration in underground environments with the objective of localizing placed targets. It prioritized target detection and target position in an absolute global frame. The evaluation criteria emphasized giving the true location of the target rather than being able to reliably guide a human or robot back to the target. Resource intensive surveying methods generated the ground truth.
@@
While recent extensions have tackled rich, multi-sensor fusion capable benchmarking [23, 9], the performance gap between SLAM benchmarking via open-loop, sensor stream replay versus task-oriented performance motivates others to evaluate SLAM and related autonomy modules through different lenses. Simulated worlds figure prominently due to reproducibility of outcomes, ease of implementation by researchers, and the availability of ground truth information [26]. Embodied AI research, such as the Habitat Navigation Challenge [27], exemplifies the necessity for a task-driven evaluation methodology. The simulation-based challenge emphasizes the need for reliable, long-term navigation capabilities. The importance of localization for autonomous vehicles motivated a simulation-based benchmarking scheme for street navigation [28] with success/failure as the main metric. SLAM latency plays a role in its robustness and accuracy, with a relationship found between these two factors for trajectory tracking tasks using SLAM pose estimates as feedback [26, 29]. Simulation-based benchmarking [26, 29] provides support for moving beyond accuracy when considering SLAM in the closed-loop. A reasonable next step would more explicitly consider navigation.
@@
An additional difference between these benchmarks and service robot deployment is that the latter often incorporates an initial mapping stage. Existing benchmarks overlook this phase by evaluating performance solely through one-time replay. SLAM maps mitigate long-term drift and ensure consistency (high precision) across repeat visits to the same locations [30, 31]. Multi-session SLAM evaluation [32] does consider map reuse with an accuracy metric for repeated runs to the same locations. However, the study design and variables may consider different performance factors and experimental methodology for LiDAR SLAM [32], and also for visual SLAM multi-map and map reuse open-loop tests [30, 31]. These works point to the need for reproducible and standardized benchmarking schemes for SLAM systems equipped with pre-built maps. Further evidence for this need is given in Fig. 1, with multi-run testing on the EuRoC dataset (see §III-C for details). The Precision vs Accuracy plots show better precision than accuracy indicating that SLAM evaluation may under-report performance. Including a mapping phase further improves performance, especially position precision. Referencing the performance metrics to robot characteristics (under $x$-axis) permits qualitative assessment of potential goal attainment success by relating precision to the visible task region at the robot’s arrival pose.
@@
This paper introduces TaskSLAM-Bench, a navigation-in-the-loop SLAM benchmark framework emphasizing task-relevant metrics: repeatability (precision) and completion. The benchmarking approach mimics real-world implementations where robot deployment may include an autonomous map building phase for improved task execution during deployment. Experiments with visual and LiDAR SLAM methods identify which methods effectively support navigation tasks. The key contributions are:
@@
  • A low-cost, easy to setup, task-driven SLAM benchmark with precision as the key metric for measuring robot consistency in reaching the same pose over multiple rounds. The method scales to large environments.
@@ II-A Preliminaries
  • Robot Pose - $(x,y,\theta)$ - Robot position and orientation. $SE(3)$ pose estimates are projected to $SE(2)$.
  • Sequence - $\mathbf{S}$ - A set of waypoints the robot is required to sequentially follow, i.e., $\mathbf{S}=\{\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{N}\}$.
  • Waypoint - $\mathbf{w}_{i}=(x_{i},y_{i},\theta_{i})$ indexed by $i$ - position and orientation (heading angle) of the robot in the 2D plane.
@@ II-B Task-Driven SLAM Benchmarking Design
Figure 2: Task-driven SLAM benchmarking design. Dashed arrows and cylinders represent optional elements specific to SLAM methods and test configurations. The $||$ symbol denotes alternative robot sensor configurations.
@@ II-B1 Task Definition
Waypoint Navigation. Waypoints represent task-relevant goal points, as exemplified in the AWS Gazebo hospital model [36] of Fig. 3. They are manually specified based on a floor plan of the environment’s free space. It is assumed that the floor plan reflects the static environment, with no dynamic obstacles or unknown areas. The robot starts at a fixed location and sequentially navigates to each waypoint. Upon reaching a waypoint, the robot’s pose is recorded for evaluation. The procedure repeats for multiple rounds, terminating if the robot fails to reach a waypoint.
@@ II-B2 Performance Metrics
Accuracy, Completeness, and Precision. Accuracy metrics such as Absolute Trajectory Error (ATE) or Relative Pose Error (RPE) typically measure SLAM performance [20]. Navigation and trajectory tracking evaluation of SLAM employ task-specific metrics [28, 26]. Here, evaluation of indoor mobile robot navigation emphasizes waypoint navigation reliability and repeatability, measured by successful goal attainment percentage (completeness) and goal pose precision. Simulated scenarios include accuracy, in the form of APE, since ground-truth is available.
@@
Task-driven Units - $(D,FoV)$. To emphasize a task-driven perspective, outcomes will also report in unit scales related to robot specific properties: $D$ the robot’s diameter, and $FoV$ the camera’s horizontal field of view.
@@ II-B3 Measuring Completeness and Precision
Waypoint precision measurement in real-world conditions may be obtained in two equivalent manners similar to virtual reality systems: outside-in or inside-out. Outside-in uses a downward facing camera from a suitable height (Fig. 4). Options include overhead/ceiling-mounted, tripod-installed, or otherwise placed to face the ground region the robot should reach. Each camera requires connection to a (low cost) computer or laptop to estimate the robot’s pose upon waypoint attainment, based on an AprilTag affixed to the top of the robot (Fig. 4). Inside-out operates as in [22] by inverting the setup: an upwards-facing camera is mounted on the robot with tags placed in the environment. AprilTags work for SLAM benchmarking [21], with the C++ detection library achieving sub-pixel accuracy ($\leq$ 0.5)[37]. Prior study [38][39] showed acceptable pose estimation accuracy for distances and viewpoints compatible with our indoor setup.
Importantly, waypoint precision does not rely on absolute references. It is assessed independently of other waypoints, with evaluation done in local coordinate systems defined by the camera and tag frames. The outside-in approach does not require global calibration of the camera poses (nor relative poses). Each camera/waypoint measurement pair operates independently. The inside-out approach approach does not require a global scan nor globally aligned tags. Waypoint attainment success is established by the tag being fully visible to the camera, otherwise waypoint attainment failed.
@@
Figure 4: Real-world experimental setup. Clock-wise from top-left: A Turtlebot2 equipped with SLAM sensors and laptop, on top of which is placed an AprilTag. Downward facing cameras (in red box) and computers/laptops placed in the environment detect and recover the robot’s pose. The robot waypoints, defined to be under the cameras, should be accurate enough that the robot will be visible when reached. During the programmed tour, the cameras will detect and estimate the robot pose when seen. Precision and Completeness follows from the estimates. The lower figure presents an example of the robot captured at two waypoints (magenta and green) for different SLAM methods. Arrows indicate orientation and black circles indicates position, over multiple runs.
@@ III Experiment Design
Effectiveness of the TaskSLAM-Bench framework will be established through a series of controlled experiments. Simulation experiments assess SLAM performance using all metrics: Accuracy, Precision, and Completeness. Plots will include dual axes with standard units and robot-relative units to link performance directly to robot characteristics and task requirements. If a robot is too far from an object, relative to its size, or headed in a bad direction, relative to its visual field of view, then a follow-up visually guided manipulation, inspection/object viewing, or other task would fail.
@@ III-A SLAM Candidates
To demonstrate the benchmark flexibility and to perform comparative analysis, several SLAM methods are tested. They include inertial and robot odometry signals when required or permitted. SLAM parameters were not tuned for optimal performance across the tests, with default settings used; they are generally effective in most scenarios and offer balanced performance. Our goal is not to identify the best SLAM method, but rather to demonstrate the functionality of the proposed SLAM benchmarking design (TaskSLAM-Bench), and to use the demonstration to assess which SLAM methods are suitable for task-driven purposes. Tested are:
@@ III-A2 Stereo Visual and Visual-Inertial
GF-GG[29] & ORB_SLAM3[30], feature-based stereo visual methods; DSOL[42], a direct odometry system; SVO-Pro[43], a semi-direct method; and MSCKF[44], a filter-based method. ORB_SLAM3 is run stereo-only as it would fail to initialize with inertial sensors, even with extra motion before testing.
@@
III-B Closed-Loop-SLAM Navigation Experiments
Testing involves both simulated and real-world environments, whose scenarios are described in Table I. In each scenario, the robot starts from a fixed initial location and navigates sequentially through the waypoints. The navigation system configuration is fixed across all experiments. Upon reaching each waypoint, the robot pauses for 5 seconds before proceeding to the next one. For an outside-in setup, the computer/laptop at this waypoint, synchronized with NTP (delay << 1 second), saves timestamped image data upon detecting the tag affixed to the robot. The robot’s poses are estimated offline and associated with waypoints by timestamp for performance evaluation.
@@ III-B2 Real World
Real-world testing was performed in an office environment. One test defined a path sequence of $\sim\!45$ meters in length and featuring six waypoints, see Fig. 4. It has short stretches in a corridor, sharp turns, and return loops. Using the outside-in approach, three overhead cameras (Fig. 4) captured the robot for success determination and precision measurement. Another test, with the inside-out approach, defined a travel sequence of about 120 meters in length with twelve waypoints. It has similar properties to the first test. Real-world sequences include nuisace factors like sensor noise, hardware delays, and low-texture areas.